OpenStudy (anonymous):

Solve the system using elimination 3x + 3y = –9 3x – 3y = 21 a. (3, –6) b. (–5, 2) c. (3, 3) d. (2, –5)

OpenStudy (anonymous):

y=-3-x 3x+9+3x=21 6x=12 x=2 6+3y=-9 3y=-15 y=-5

OpenStudy (anonymous):

so the answer is d.

OpenStudy (anonymous):

d. (2, –5) Subtract: -> -6y = 30 => y = - 30/6 = -5 => x = 7 + y = 2